OpenStudy (anonymous):

Show all work to solve 2 over x plus 1 over 9 equals 1 over 3. Use complete sentences to explain how to solve this equation.

OpenStudy (anonymous):

(2/x) + (1/9) = (1/3)... is that right? Change 1/3 to 3/9 so you have the two fractions with the same denominator. That leaves: (2/x) + (1/9) = (3/9) Then subtract the 1/9 from both sides: (2/x) = (2/9) Time both sides by x: (2) = (2/9)x Divide by (2/9) on both sides: (2)/(2/9) = x Simplify: 9 = x